### v12

# Hungarian Research Network (HUN-REN) Laboratory on Engineering and Management Intelligence Research Group of Operations Research and Decision Systems

Abstract
Several sports tournaments contain a round-robin group stage where the teams are assigned to groups subject to some constraints. Hence, the organisers usually use a computer-assisted random draw to avoid any dead end, a situation when the teams still to be drawn cannot be assigned to the remaining empty slots. This procedure is known to be unfair: the feasible allocations are not equally likely. We quantify the unfairness of the 2018 FIFA World Cup draw and evaluate its effect on the probability of qualification for the knockout stage for each national team. The official draw order of Pot 1, Pot 2, Pot 3, Pot 4 turns out to be a significantly better option than the 23 other draw orders with respect to the unwanted distortions. Nonetheless, the non-uniform draw distorts the probability of qualification by more than one percentage point for two countries. Our results call attention to the non-negligible role of draw order and make it possible for policymakers to decide whether the use of fairer draw mechanisms is justified.
Keywords: fairness; FIFA World Cup; group draw; permutation; simulation
MSC class: 05A05, 68W40, 90-10, 91B14
JEL classification number: C44, C63, Z20
Any one who considers arithmetical methods of producing random digits is, of course, in a state of sin.111 Source: von Neumann, (1951, p. 36).
(John von Neumann: Various techniques used in connection with random digits)
[…]
1 Introduction
The mechanism design literature usually focuses on theoretical requirements like efficiency, fairness, and incentive compatibility (Abdulkadiroğlu and Sönmez,, 2003; Roth et al.,, 2004; Csató,, 2021). On the other hand, institutions—like governing bodies in major sports—often emphasise simplicity and transparency, which calls for a comprehensive review of how the procedures that exist in the real world perform with respect to the above properties.
Many sports tournaments involve a group stage where the teams are assigned to groups subject to some constraints (Kobierecki,, 2022; Laliena and López,, 2019). These constraints are imposed “to issue a schedule that is fair for the participating teams, fulfils the expectations of commercial partners and ensures with a high degree of probability that the fixture can take place as scheduled(UEFA, 2020a, ). Examples include the FIBA Basketball World Cup (FIBA,, 2023), the FIFA World Cup (FIFA,, 2022), the European Qualifiers for the FIFA World Cup (UEFA, 2020a, ), the UEFA Euro qualifying (UEFA,, 2022), and the UEFA Nations League (UEFA,, 2024). Draw restrictions are also used in club-level tournaments such as the UEFA Champions League and the UEFA Europa League (Csató, 2022a, ).
Until 2014, the FIFA World Cup draw divided the teams into pots mainly according to their geographic area, which caused serious unfairness in certain years such as in 1990 (Jones,, 1990), 2006 (Rathgeber and Rathgeber,, 2007), and 2014 (Guyon,, 2015). In particular, the top teams had different chances of being placed in a group with difficult opponents. Consequently, Guyon, (2014) has recommended three options to create balanced and geographically diverse groups, which has inspired FIFA to change the draw procedure used in the 2018 and 2022 FIFA World Cups (Guyon,, 2018; Csató,, 2023).
However, this mechanism, based on a computer-assisted random draw to avoid any dead end (a situation when the teams still to be drawn cannot be assigned to the remaining empty slots), is not uniformly distributed: the valid assignments are not equally likely to occur (Roberts and Rosenthal,, 2023). Hence, Roberts and Rosenthal, (2023) propose two unbiased mechanisms that use balls and bowls, making them suitable for a televised draw. On the other hand, they require random simulations, which might threaten transparency.
Therefore, the organiser faces a dilemma: retain the existing but biased method or switch to a correct but less transparent mechanism. In order to understand this trade-off and choose the better option, it is inevitable to explore the extent of the bias, as well as its potential sporting effects. While these issues have been recently analysed in the UEFA Champions League (Boczoń and Wilson,, 2023), they have not been addressed in the case of the FIFA World Cup, the most prominent football tournament around the world.
The current work aims to fill this research gap. In particular, the 2018 FIFA World Cup has been chosen as a case study since the 2022 FIFA World Cup draw has contained an inherent bias due to the uncertainty in the set of teams qualified (Csató,, 2023), and the teams qualifying for the 2026 FIFA World Cup are naturally unknown at the time of writing.
We also call attention to the role of draw order: the order of the pots in the group draw turns out to have a non-negligible sporting effect. This is a somewhat surprising finding because the fairness of the UEFA Champions League Round of 16 draw essentially does not depend on whether the group winners or the runners-up are drawn first (Boczoń and Wilson,, 2023; Klößner and Becker,, 2013).
Our topic is hugely relevant to governing bodies in sports. In an ideal environment, the rules governing a sport prevent any player or team from gaining an unfair advantage (Csató,, 2021; Devriesere et al.,, 2024). But the sequential draw method of FIFA may threaten fairness, while the alternative solutions remain more complex and less transparent (Roberts and Rosenthal,, 2023). Therefore, policymakers could maintain the balance between fairness and integrity only by investigating different fairness indicators (Scelles et al.,, 2024).
The paper starts with presenting a concise review of related literature, followed by describing the theoretical background of the FIFA World Cup group draw. The numerical results for the 2018 FIFA World Cup draw are discussed in two parts. First, the departure of the draw procedure from a uniformly distributed random choice among all feasible allocations is quantified for all the 24 possible draw orders of the four pots. This is important because the modification of the draw order (relabelling of the pots) does not require any reform in the existing principles of the draw. Second, the bias of the draw procedure is evaluated with respect to the probability of qualification for the knockout stage. That is essential since the ultimate price to pay for the sake of public interest and transparency is the distortion of the final outcome: if the effects on the chances of the teams remain marginal and insignificant, then there is no need to choose a more complex and less transparent draw procedure.
2 Related literature
Several scientific works analyse the FIFA World Cup draw. Before the 2018 edition, the host nation and the strongest teams were assigned to different groups, while the remaining teams were drawn randomly with maximising geographic separation: countries from the same continent (except for Europe) could not have played in the same group and at most two European teams could have been in any group.
In the case of the 1990 FIFA World Cup, Jones, (1990) shows that the draw was not mathematically fair. For example, West Germany would have been up against a South American team with a probability of 4/5 instead of 1/2—as it should have been—due to the incorrect consideration of the constraints. Similarly, the host Germany was likely to play in a difficult group in 2006, but other seeded teams, such as Italy, were not (Rathgeber and Rathgeber,, 2007).
Guyon, (2015) identifies severe shortcomings of the mechanism used for the 2014 FIFA World Cup draw such as imbalance (the eight groups are at different competitive levels), unfairness (certain teams have a greater chance of ending up in a tough group), and non-uniform distribution (the feasible allocations are not equally likely). Indeed, there has been a substantial competitive imbalance between the historical FIFA World Cup groups (Lapré and Palazzolo,, 2023).
Guyon, (2014) presents alternative proposals to retain the practicalities of the FIFA World Cup draw but improve its outcome. One of them can be compared to the flawed FIFA rule at https://www.nytimes.com/interactive/2014/06/03/upshot/world-cup-draw-simulation.html.
Laliena and López, (2019) develop two uniformly distributed designs for group draw with geographical restrictions that produce groups having similar or equal competitive levels. Cea et al., (2020) analyse the deficiencies of the 2014 FIFA World Cup draw and provide a mixed integer linear programming model to create the groups. The suggested method takes into account draw restrictions and aims to balance “quality” across the groups.
Roberts and Rosenthal, (2023) consider the challenge of finding a group draw mechanism that follows the uniform distribution over all valid assignments but is also entertaining, practical, and transparent. The authors suggest two procedures for achieving this aim by using balls and bowls in a way, which is suitable for a nice television show—but, in contrast to the proposals of Guyon, (2014), they use computer draws at some stage which may threaten transparency. Both algorithms can be tried interactively at http://probability.ca/fdraw/.
Other studies deal with the UEFA Champions League where a constrained draw mechanism has been used in the Round of 16 between the 2003/04 and 2023/24 seasons. However, this method is distinct from the FIFA World Cup draw procedure (Csató,, 2024). Wallace and Haigh, (2013) verify that the possible assignments are not equally likely and highlight the connection of the UEFA draw mechanism to Hall’s marriage theorem, see also Haigh, (2019, Section 3.6). Kiesl, (2013) computes the bias in the 2012/13 season and outlines some fair—but uninteresting to watch—methods. According to Klößner and Becker, (2013), the draw system inherently implies different probabilities for certain assignments, which are translated into more than ten thousand Euros in expected revenue due to the substantial amount of prize money. Finally, Boczoń and Wilson, (2023) reveal how the UEFA draw procedure affects expected assignments and address the normative question of whether a fairer randomisation mechanism exists. The current design is verified to come quantitatively close to a constrained best in fairness terms.
3 Theoretical background
A permutation of an ordered set is a rearrangement of its elements. In the FIFA World Cup draw, the initial order of the teams is provided by a random draw. In an unrestricted group draw, the teams can be assigned to the groups in this order. However, in the presence of draw conditions, it is not obvious to find the permutation of the teams that corresponds to the feasible allocation implied by the draw procedure of FIFA.
This mechanism is defined as follows: “when a draw condition applies or is anticipated to apply, the team drawn is allocated to the first available group in alphabetical order(UEFA, 2020a, ). In other words, the team drawn is assigned to the first empty slot except if all permutations of the remaining teams violate at least one draw condition.
Group Team assignment: the first 12 permutations
1 2 3 4 5 6 7 8 9 10 11 12
A
B
C
D
Group Team assignment: the last 12 permutations
13 14 15 16 17 18 19 20 21 22 23 24
A
B
C
D
  • The order of the teams according to the random draw is $T1$$T2$$T3$$T4$.
  • The symbols , , , represent teams $T1$$T4$ in Example 1, respectively.
Figure 1: The sequence of permutations implied by the draw procedure of FIFA, $n=4$
Example 1.
Assume that there are $k=4$ groups A–D and $n=4$ teams $T1$$T4$ drawn sequentially from a pot. The order of permutations to be checked according to the draw mechanism is shown in Figure 1. Team $T1$ is assigned to group A in the first six permutations because it can be placed in another group only if either group A is unavailable for team $T1$ or teams $T2$$T4$ cannot be assigned to groups B–D.
Let us consider two illustrative cases, where the restrictions are implied by the assignment of teams drawn from the previous pot(s):
  • If team $T1$ cannot be placed in group A and team $T3$ cannot be placed in group C, then the first six permutations are unacceptable due to the first constraint, and permutation 7 is skipped because of the second condition. The outcome of the draw procedure is permutation 8 ($T2$, $T1$, $T4$, $T3$).
  • If teams $T2$$T4$ cannot be placed in group C and team $T2$ cannot be placed in group A, then the first 12 permutations are unacceptable due to the first constraint, and the next two are skipped because of the second condition. The outcome of the draw procedure is permutation 15 ($T3$, $T2$, $T1$, $T4$).
Generating all permutations of a sequence of values in a specific order is a famous problem in computer science (Sedgewick,, 1977). The classic algorithm of lexicographic ordering goes back to Narayaṇa Paṇḍita, an Indian mathematician from the 14th century (Knuth,, 2005). The ordering corresponding to the FIFA World Cup draw procedure is called representation via swaps (Arndt,, 2010) and has been presented first in Myrvold and Ruskey, (2001) according to our knowledge. Arndt, (2010, Figure 10.1-E) contains the same order of permutations as Figure 1.
[…]
Figure 2: A backtracking algorithm for restricted group draw that finds
the feasible allocation corresponding to a given order of the teams
The description of the 2018 FIFA World Cup draw (FIFA,, 2017) does not give an exact algorithm to obtain the implied feasible allocation of the teams into groups for a given random order. The scheme of an appropriate computer program is presented in Figure 2. Pot $j$, from which team $i$ is drawn, is not a variable, however, the number of available groups can be different for each pot. The algorithm is based on backtracking: if the remaining teams cannot be assigned to the empty group slots in any order such that all restrictions are satisfied, then the last team is placed in the next available group in alphabetical order. This process is repeated until the associated feasible allocation is obtained or the non-existence of a valid assignment is verified.
Backtracking can be familiar from the problem of scheduling round-robin tournaments, where an unlucky assignment of games to slots can result in a schedule that cannot be completed (Rosa and Wallis,, 1982; Schaerf,, 1999). Backtracking is also widely used to solve puzzles such as the eight queens puzzle, crosswords, or Sudoku.
4 The (un)fairness of the 2018 FIFA World Cup draw
For the 2018 FIFA World Cup draw, pots were constructed based on the October 2017 FIFA World Ranking such that Pot $j$ contained the teams ranked between $18(j-1)+$ 1 and $8j$. The only exception was the assignment of the host Russia to Pot 1 despite being the lowest-ranked among all participants.
The draw sequence started with Pot 1 and ended with Pot 4. Each pot was emptied before the next was drawn and some draw conditions were applied (FIFA,, 2017):
  • Russia was automatically placed in Group A.
  • No group could have more than one team from any continental confederation except for UEFA (AFC, CAF, CONMEBOL, CONCACAF).
  • Each group should have contained at least one but no more than two European teams.
Table 1: Seeding pots in the 2018 FIFA World Cup
Country Confederation Elo Country Confederation Elo
Pot 1 Pot 2
1 Russia UEFA 1678 90 Spain UEFA 2044
2 Germany UEFA 2077 10 Peru CONMEBOL 1916
3 Brazil CONMEBOL 2141 11 Switzerland UEFA 1889
4 Portugal UEFA 1969 12 England UEFA 1948
5 Argentina CONMEBOL 1985 13 Colombia CONMEBOL 1927
6 Belgium UEFA 1937 14 Mexico CONCACAF 1850
7 Poland UEFA 1831 15 Uruguay CONMEBOL 1893
8 France UEFA 1986 16 Croatia UEFA 1853
Pot 3 Pot 4
17 Denmark UEFA 1856 25 Serbia UEFA 1777
18 Iceland UEFA 1764 26 Nigeria CAF 1684
19 Costa Rica CONCACAF 1743 27 Australia AFC 1741
20 Sweden UEFA 1795 28 Japan AFC 1684
21 Tunisia CAF 1655 29 Morocco CAF 1733
22 Egypt CAF 1643 30 Panama CONCACAF 1658
23 Senegal CAF 1749 31 South Korea AFC 1713
24 Iran AFC 1790 32 Saudi Arabia AFC 1586
  • The number before each country indicates its rank among the FIFA World Cup participants according to the October 2017 FIFA World Ranking, except for the host Russia, which automatically occupies the first position.
  • The column Elo shows the strength of the teams according to the World Football Elo Ratings on 13 June 2018, see https://www.international-football.net/elo-ratings-table?year=2018&month=06&day=13&confed=&. The 2018 FIFA World Cup started on 14 June 2018. This measure quantifies the strengths of the teams in our simulation.
The composition of the pots is shown in Table 1.
4.1 The effect of draw order
According to Roberts and Rosenthal, (2023, Section 2), the FIFA World Cup draw procedure is unfair since some feasible allocations might occur with a higher probability. In addition, the pre-assignment of Russia to Group A introduces a powerful bias because the draw mechanism is not independent of group labels. Russia has a 12.5% probability of playing against an arbitrarily chosen country from Pot 2 because no draw constraints can apply. However, since there are one CONCACAF, three CONMEBOL, and four UEFA members in Pot 2, the two CONMEBOL teams from Pot 1 (Brazil and Argentina) play against a given European team from Pot 2 with a probability of 0.2 since they have five possible opponents from Pot 2. The remaining five UEFA teams in Pot 1 are identical concerning the draw constraints, thus, they have a chance of $(100-12.5)/5=17.5$% to be assigned to the same group as a given South American team from Pot 2. Analogously, the probability that two given European teams from Pots 1 and 2 play against each other equals $(100-2\times 20-12.5)/5=9.5$%.
However, an appropriate relabelling of the pots may bring the FIFA World Cup draw closer to the uniform distribution. Therefore, all possible orders of the pots are examined such that the pre-assignment of Russia to Group A is retained. Since the teams can be drawn in $7!\times\left(8!\right)^{3}\approx 3.3\times 10^{17}$ different orders, it is almost impossible to derive exact theoretical results. Consequently, the 25 draw mechanisms—the draw procedure of FIFA with the 24 possible orders of the pots and the uniformly distributed rejection mechanism (Roberts and Rosenthal,, 2023)—are analysed on the basis of 1 million randomly generated draw orders, analogous to Roberts and Rosenthal, (2023).
The distortions compared to the rejection mechanism $U$ are quantified through the difference of the probabilities that two given teams are placed in the same group. The bias $\Delta_{ij}$ of mechanism $M$ for teams $i$ and $j$ is suggested to be
$\Delta_{ij}=p_{ij}^{M}-p_{ij}^{U},$
where $p_{ij}^{M}$ and $p_{ij}^{U}$ are the probabilities that teams $i$ and $j$ are assigned to the same group by a mechanism $M$ and the uniform draw, respectively. When the biases are aggregated for the teams, we will always sum their absolute value $\lvert\Delta_{ij}\rvert$.
Table 2: The deviations of different draw mechanisms in the 2018 FIFA World Cup
Draw order Minimal nonzero probability Maximal probability Maximum of positive ($+$) biases Maximum of negative (--) biases Sum of absolute biases
1-2-3-4 3.26% 32.98% 10.29% (19) 4.68% (7) 3.007 (1)
1-2-4-3 1.84% 29.62% 09.97% (18) 4.18% (1) 3.122 (3)
1-3-2-4 1.18% 28.36% 08.46% (15) 5.17% (11) 3.867 (14)
1-3-4-2 2.86% 28.41% 03.86% (4) 5.19% (13) 3.786 (10)
1-4-2-3 1.05% 29.55% 07.35% (14) 6.33% (19) 3.364 (6)
1-4-3-2 2.12% 28.17% 03.21% (1) 6.33% (18) 3.756 (8)
2-1-3-4 3.26% 33.29% 10.60% (20) 4.67% (6) 3.104 (2)
2-1-4-3 1.75% 29.64% 09.96% (17) 4.23% (2) 3.206 (4)
2-3-1-4 1.00% 28.44% 06.08% (9) 5.69% (16) 5.174 (22)
2-3-4-1 1.67% 28.62% 06.02% (8) 4.87% (8) 4.825 (21)
2-4-1-3 1.11% 29.43% 06.84% (11) 7.45% (23) 3.810 (11)
2-4-3-1 1.48% 28.47% 06.74% (10) 7.53% (24) 4.217 (16)
3-1-2-4 1.18% 28.36% 08.47% (16) 5.17% (12) 3.844 (12)
3-1-4-2 2.85% 28.50% 03.81% (3) 5.17% (10) 3.775 (9)
3-2-1-4 0.95% 28.64% 05.71% (7) 5.29% (15) 4.628 (18)
3-2-4-1 1.37% 28.68% 05.48% (6) 5.07% (9) 4.397 (17)
3-4-1-2 4.13% 39.07% 16.38% (24) 4.52% (4) 4.653 (19)
3-4-2-1 5.72% 35.46% 12.77% (22) 5.26% (14) 5.619 (23)
4-1-2-3 1.05% 29.59% 07.33% (13) 6.27% (17) 3.341 (5)
4-1-3-2 2.12% 28.21% 03.28% (2) 6.38% (20) 3.859 (13)
4-2-1-3 1.09% 29.63% 06.94% (12) 6.83% (22) 3.503 (7)
4-2-3-1 1.35% 28.53% 05.24% (5) 6.76% (21) 4.138 (15)
4-3-1-2 4.11% 37.96% 15.27% (23) 4.24% (3) 4.656 (20)
4-3-2-1 5.47% 34.45% 11.76% (21) 4.67% (5) 5.671 (24)
Uniform 1.56% 28.74%
  • The numbers in parenthesis indicate the ranks of FIFA draw procedure with the draw order in the row according to the measure of deviation in the column.
The most extreme and aggregated distortions are presented in Table 2. For example, the probability that Denmark (17) (or the equivalent team of Sweden (18) or Iceland (20)) and Serbia (25) play in the same group is more than doubled by the FIFA draw mechanism with the traditional draw order 1-2-3-4. On the other hand, the likelihood of assigning Russia (1) and Serbia to the same group is decreased by 4.68 percentage points, while the likelihood of assigning Mexico (14) and Serbia to the same group is increased by 10.29 percentage points.
The standard error of $\Delta_{ij}$ equals
[…]
The last column of Table 2 compares the 24 draw orders by adding the values of $\Delta_{ij}$ for the 365 allowed country pairs. The bounds of the 99% confidence intervals are given by
[…]
Hence, the official draw order 1-2-3-4 is not only optimal but it is significantly fairer than any other draw order. Furthermore, Table 2 uncovers that there are some low-performing draw orders, dominated by another draw order, such as 3-4-2-1 and 4-3-2-1.
The distortions are also worth studying from the perspective of the countries by taking the average of absolute or squared biases for the pairs that involve the given national team. For the sake of simplicity, four draw orders will be analysed in detail: the official 1-2-3-4, one that improves both the minimal nonzero and maximal probabilities with a slight change (1-2-4-3), one that is less biased for Mexico (1-3-2-4), and a fundamentally different order (4-2-3-1).
[…]
Figure 3: The average biases of different draw procedures
for the national teams in the 2018 FIFA World Cup
Figure 3 compares the means of the biases for the 32 countries. The official draw procedure is far from the fair rejection mechanism in the case of Russia (1), Brazil (3), Argentina (5), Mexico (14), and Serbia (25). Note that Brazil and Argentina are interchangeable in the draw, thus, the deviations for these two countries are different only due to the (seemingly low) sampling error. Reversing the order of Pots 3 and 4 reduces unfairness for Russia and Serbia but it does not treat the problem for Mexico. Contrarily, the deviation for Costa Rica (19) strongly increases with the draw order 1-2-4-3. Drawing Pot 3 immediately after the best teams in Pot 1 implies a relatively low distortion in the case of Russia, and, especially, Mexico but the situation of the African (21, 22, 23) and European (17, 18, 20) teams in Pot 3 becomes less fair compared to the draw order 1-2-4-3. Finally, the draw order 4-2-3-1 is unfair with respect to all countries in Pot 2 except for Mexico.
4.2 The consequences of non-uniform distribution
In the 2018 FIFA World Cup, the top two teams from each group have advanced to the Round of 16. Therefore, the distortions of the draw procedure are important primarily if they affect the probability of qualification for the knockout stage. To that end, the simulation methodology of Football rankings, (2020) is used. This models the number of goals scored in a match by Poisson distribution: the expected number of goals is a quartic polynomial of win expectancy as estimated by a least squares regression based on more than 29 thousand home-away games and almost 10 thousand games played on neutral ground between national football teams (Football rankings,, 2020). Win expectancy depends on the strengths of the teams according to a well-established metric (Lasek et al.,, 2013; Gásquez and Royuela,, 2016), the World Football Elo ratings (http://eloratings.net/about), and the field of the match, which is neutral except for Russia, the host.
In particular, the probability that team $i$ scores $k$ goals against team $j$ equals
$P_{ij}(k)=\frac{\left(\lambda_{ij}^{(f)}\right)^{k}\exp\left(-\lambda_{ij}^{(f)}\right)}{k!},$
where $\lambda_{ij}^{(f)}$ is the expected number of goals scored by team $i$ against team $j$ if the match is played on field $f$ (home: $f=h$; away: $f=a$; neutral: $f=n$).
World Football Elo ratings determine the win expectancy $W_{ij}$ of team $i$ against team $j$ as
$W_{ij}=\frac{1}{1+10^{-(E_{i}-E_{j})/400}},$
where $E_{i}$ and $E_{j}$ are the Elo ratings of the two teams, respectively. The rating of the home team (here, Russia) is increased by $100$ to reflect home advantage.
Football rankings, (2020) estimates how $\lambda_{ij}^{(f)}$ depends on $W_{ij}$ by least squares regressions with a regime change at $W_{ij}=0.9$ due to the excessive number of goals scored in unbalanced matches. Most games of the FIFA World Cup are played on neutral field when
$\lambda_{ij}^{(n)}=\left\{\begin{array}[]{ll}3.90388\cdot W_{ij}^{4}-0.58486\cdot W_{ij}^{3}\\ -2.98315\cdot W_{ij}^{2}+3.13160\cdot W_{ij}+0.33193&\textrm{if }W_{ij}\leq 0.9\\ \\ 308097.45501\cdot(W_{ij}-0.9)^{4}-42803.04696\cdot(W_{ij}-0.9)^{3}&\\ +2116.35304\cdot(W_{ij}-0.9)^{2}-9.61869\cdot(W_{ij}-0.9)+2.86899&\textrm{if }W_{ij}>0.9.\end{array}\right.$
In Group A, there are three home-away matches played by Russia and another team. Then the expected number of goals scored by the host Russia (denoted by $R$) equals
$\lambda_{Rj}^{(h)}=\left\{\begin{array}[]{ll}-5.42301\cdot W_{Rj}^{4}+15.49728\cdot W_{Rj}^{3}\\ -12.6499\cdot W_{Rj}^{2}+5.36198\cdot W_{Rj}+0.22862&\textrm{if }W_{Rj}\leq 0.9\\ \\ 231098.16153\cdot(W_{Rj}-0.9)^{4}-30953.10199\cdot(W_{Rj}-0.9)^{3}&\\ +1347.51495\cdot(W_{Rj}-0.9)^{2}-1.63074\cdot(W_{Rj}-0.9)+2.54747&\textrm{if }W_{Rj}>0.9,\end{array}\right.$
and the expected number of goals scored by the away team $j$ is
$\lambda_{Rj}^{(a)}=\left\{\begin{array}[]{ll}90173.57949\cdot(W_{Rj}-0.1)^{4}+10064.38612\cdot(W_{Rj}-0.1)^{3}\\ +218.6628\cdot(W_{Rj}-0.1)^{2}-11.06198\cdot(W_{Rj}-0.1)+2.28291&\textrm{if }W_{Rj}<0.1\\ \\ -1.25010\cdot W_{Rj}^{4}-1.99984\cdot W_{Rj}^{3}&\\ +6.54946\cdot W_{Rj}^{2}-5.83979\cdot W_{Rj}+2.80352&\textrm{if }W_{Rj}\geq 0.1.\end{array}\right.$
The same simulation methodology has recently been applied in some studies on tournament design (Csató, 2022b, ; Csató,, 2023; Stronka,, 2024). Note that even if this Elo-based approach is not necessarily the best available simulation model, it is only used to compare the outcomes under the uniform and non-uniform draw. Thus, our assumptions seem to be adequate for comparative purposes as emphasised by Appleton, (1995).
22-2- 211-1- 100111155551010101015151515202020202525252530303030Change in percentage pointsRank of the national team in the ranking used for the drawChange in the probability of winning the groupOfficial (draw order 1-2-3-4)      Draw order 1-2-4-3Draw order 1-3-2-4                  Draw order 4-2-3-111-1- 10.50.5-0.5- 0.5000.50.50.50.5111155551010101015151515202020202525252530303030Change in percentage pointsRank of the national team in the ranking used for the drawChange in the probability of being the runner-up
Figure 4: The effect of different draw procedures on being the
group winner and the runner-up in the 2018 FIFA World Cup
Figure 4 shows how the FIFA World Cup draw mechanism distorts the chances of winning the group and being the runner-up. The official rule (draw order 1-2-3-4) increases the probability of winning the group by more than 1 percentage point for the four UEFA teams in Pot 2 (9, 11, 12, 16) mostly at the expense of Mexico (14) and Serbia (25). Fortunately, these effects are somewhat mitigated by taking the likelihood of obtaining the second position into account. While draw order 1-2-4-3 does not differ much from the traditional order of 1-2-3-4, draw order 1-3-2-4 strongly reduces the distortions in the case of the above countries except for Serbia. Among the four draw orders, 4-2-3-1 is the less favourable for the weakest European team in Pot 2 (16) and the three UEFA members in Pot 3 (17, 18, 20).
1.51.5-1.5- 1.511-1- 10.50.5-0.5- 0.5000.50.50.50.511111.51.51.51.5RussiaGermanyBrazilPortugalArgentinaBelgiumPolandFranceSpainPeruSwitzerlandEnglandColombiaMexicoUruguayCroatiaDenmarkIcelandCosta RicaSwedenTunisiaEgyptSenegalIranSerbiaNigeriaAustraliaJapanMoroccoPanamaSouth KoreaSaudi ArabiaChange in percentage pointsAbsolute changes in the probability of qualificationOfficial (draw order 1-2-3-4)      Draw order 1-2-4-3Draw order 1-3-2-4                  Draw order 4-2-3-1
Figure 5: The absolute effect of different draw procedures on the probability
of qualification for the knockout stage in the 2018 FIFA World Cup
Table 3: Summary statistics on the effects of four
draw procedures in the 2018 FIFA World Cup
Draw order 1-2-3-4 1-2-4-3 1-3-2-4 4-2-3-1
Change GW RU Q GW RU Q GW RU Q GW RU Q
$>1\%$ 4 0 1 3 0 0 0 0 1 1 0 1
$0.5\%$$1\%$ 2 2 2 3 2 2 3 0 2 1 2 5
$0.25\%$$0.5\%$ 1 4 8 1 5 4 2 4 5 7 6 8
$-0.25\%$$0.25\%$ 12 20 11 13 18 20 20 24 14 17 17 8
$-0.5\%$$-0.25\%$ 6 0 5 7 1 4 6 4 7 1 4 2
$-1\%$$-0.5\%$ 6 6 4 4 6 1 0 0 2 4 3 4
$<-1\%$ 1 0 1 1 0 1 1 0 1 1 0 4
  • Column GW shows the number of teams for which the change in the probability of winning the group is within the given interval.
  • Column RU shows the number of teams for which the change in the probability of being the runner-up is within the given interval.
  • Column Q shows the number of teams for which the change in the probability of qualification is within the given interval.
The probabilities of being the group winner or the runner-up are aggregated in Figure 5. Draw order 1-2-4-3 is somewhat better than the official 1-2-3-4: despite the increased impact on Costa Rica, it is less unfair for Russia, Croatia, Denmark, Iceland, and Sweden. Draw order 1-3-2-4 can be chosen if the bias for Mexico should be reduced, however, that is achieved at the expense of Serbia. Draw order 4-2-3-1 is not worth implementing because of the high distortions for several countries. These observations are reinforced by the summary statistics provided in Table 3.
44-4- 433-3- 322-2- 211-1- 1001111222233334444RussiaGermanyBrazilPortugalArgentinaBelgiumPolandFranceSpainPeruSwitzerlandEnglandColombiaMexicoUruguayCroatiaDenmarkIcelandCosta RicaSwedenTunisiaEgyptSenegalIranSerbiaNigeriaAustraliaJapanMoroccoPanamaSouth KoreaSaudi ArabiaChange in %Relative changes in the probability of qualificationOfficial (draw order 1-2-3-4)      Draw order 1-2-4-3Draw order 1-3-2-4                  Draw order 4-2-3-1
Figure 6: The relative effect of different draw procedures on the probability
of qualification for the knockout stage in the 2018 FIFA World Cup
Table 4: Summary statistics on the relative effects of
four draw procedures in the 2018 FIFA World Cup
Change 1-2-3-4 1-2-4-3 1-3-2-4 4-2-3-1
$>3\%$ 0 0 0 2
$2\%$$3\%$ 3 1 2 2
$1\%$$2\%$ 7 4 3 3
$-1\%$$1\%$ 16 24 21 17
$-2\%$$-1\%$ 4 2 4 4
$-3\%$$-2\%$ 2 1 1 3
$<-3\%$ 0 0 1 1
  • The table reports the number of national teams for which the change in the probability of qualification by the given draw order is within the appropriate interval.
Finally, Figure 6 uncovers the relative changes in the probability of qualification for the knockout stage as the same absolute distortions can be more costly for the weaker teams. In this respect, draw order 1-2-4-3 can be a reasonable alternative to the official 1-2-3-4 as it is less biased for almost all teams except for Costa Rica and Nigeria as shown by Table 4. The other two orders are especially unfavourable for Serbia and the three UEFA (draw order 1-2-4-3) or the three CAF (draw order 4-2-3-1) teams in Pot 3.
Based on the arguments above, the official draw order of Pot 1, Pot 2, Pot 3, Pot 4 has been a lucky choice for the 2018 FIFA World Cup draw. But a change in the continental allocation of the teams playing in the FIFA World Cup may lead to a highly unfair draw in the future if the current draw procedure is applied.
5 Discussion
Our paper has analysed the unfairness of the 2018 FIFA World Cup draw. First, the connection of the standard constrained group draw procedure—used by the FIFA and UEFA—to permutation generation has been presented. Second, we have examined how this procedure departs from a random draw among all feasible allocations and considered all alternatives by relabelling the pots. The official draw order (Pot 1, Pot 2, Pot 3, Pot 4) has turned out to be an optimal choice among the 24 alternative rules with respect to natural measures of fairness.
We think there is at least some scope to modify the current draw order because the draw of certain sports tournaments has not been started with the strongest teams:
  • The lowest-ranked teams were drawn first in the 2020/21 (UEFA, 2020b, ) and 2022/23 (UEFA,, 2021) UEFA Nations League, as well as in the 2019 World Men’s Handball Championship (IHF,, 2018);
  • The runners-up were drawn first in the UEFA Champions League Round of 16 (Boczoń and Wilson,, 2023; Klößner and Becker,, 2013), and the unseeded teams were drawn first in the UEFA Europa League Round of 32 (Csató, 2022a, );
  • The unusual draw order of Pot 4, Pot 3, Pot 1, Pot 2 was followed in the draw of the 2021 World Men’s Handball Championship (IHF,, 2020).
According to our findings, the non-uniform draw distorts the probability of qualification for eight (two) countries by more than 0.5 (1) percentage points (see Table 3). Although the bias has not exceeded 1.5 percentage points for any national team in the case of the 2018 FIFA World Cup, it has exceeded 2% for Russia, Mexico, Iceland, Tunisia, and Egypt in relative terms. To conclude, it is the responsibility of policymakers to decide whether these values justify the consideration of fairer draw mechanisms such as the algorithms proposed by Guyon, (2014) and Roberts and Rosenthal, (2023).
Acknowledgements
This paper could not have been written without my father (also called László Csató), who has primarily coded the simulations in Python.
We are grateful to Julien Guyon for inspiration, and to Julien Guyon, Jeffrey S. Rosenthal, and Lajos Rónyai for useful advice.
Eleven colleagues and anonymous reviewers provided valuable remarks and suggestions on earlier drafts.
We are indebted to the Wikipedia community for summarising important details of the sports competition discussed in the paper.
The research was supported by the National Research, Development and Innovation Office under Grant FK 145838.